# Online Housing Market

### v1

@@ 7 Top Trading Cycle Based Algorithms
Let me first show that Algorithm 6, denoted $\mathcal{A}$ in this proof, is online when the partition function at use is $SOC$. Consider an instance $I$ and an agent $i$. Let $I^{\prime}$ denote the restricted instance $I_{<d_{i}}$. I need to show that $\mathcal{A}_{i}(I)=\mathcal{A}_{i}(I^{\prime})$. Since $a_{i}<d_{j}$ holds for any agent $j$ of $I^{\prime}$, it follows that $I_{<d_{j}}=I^{\prime}_{<d_{j}}$. This implies that $\mathcal{P}_{j}(I_{<d_{j}})=\mathcal{P}_{j}(I^{\prime}_{<d_{j}})$ holds for any agent $j$ in $I^{\prime}$. Therefore, the partitions of agents in $I^{\prime}$ are the same in $\mathcal{P}(I)$ and $\mathcal{P}(I^{\prime})$, and Algorithm 6, applied to both $I$ and $I^{\prime}$, executes the same iterations up to the departure time $d_{i}$ of agent $i$. This implies that $\mathcal{A}_{i}(I)=\mathcal{A}_{i}(I^{\prime})$.
Let me now show that it is $WIC$. It is well known that Algorithm 5 is incentive compatible. During Algorithm 6, multiple instances of Algorithm 5 are applied, but each agent participates in exactly a single one of them. Furthermore, I assumed that the partition function used to select the agents participating in each instance of Algorithm 5 does not depend on the agents’ preferences. Therefore, no agent has an incentive to misreport her preferences, and Algorithm 6 is $WIC$. ∎
@@ 7.1 The Departing Agent Excluded Partition
I explore now if a stronger notion of incentive compatibility can be achieved with particular partition functions. Let me introduce the departing agent excluded partition $\gamma$, whose construction is described in Algorithm 7. Intuitively, every time that an unpartitioned agent leaves the market, two new subsets are added to the partition. The first is a singleton containing the agent leaving the market, and the second is the remaining unpartitioned agents that are present in the market. The following proposition shows that this partition function is static online compatible.
@@
The departing agent excluded partition $\gamma$ is $SOC$
@@
I need to show that for any instance $I$ and for any agent $i$, $\gamma_{i}(I)=\gamma_{i}(I_{<d_{i}})$ holds. Let $i^{\prime}$ denote the rank of agent $i$ in the ascending departure permutation $\delta$, i.e., such that $\delta(i^{\prime})=i$ holds. Note that for any position $j\leq i^{\prime}$, agents in $N_{<d_{\delta(j)}}$ are all contained in $I_{<d_{i}}$ since $d_{i}\geq d_{\delta(j)}$. Therefore, the first $i^{\prime}$ iterations of the “for” loop of Algorithm 7 are the same for $I$ and $I_{<d_{i}}$. Consequently, the subset of agents containing agent $i$ in the departing agent excluded partition $\gamma$ is the same for both instances $I$ and $I_{<d_{i}}$. ∎
@@
Consider the instance $I$ depicted in the left part of Figure 4, and the three allocations described in the right part of the same figure. It is easy to verify that Algorithm 1 using either the ascending departure permutation $\delta$ or the ascending arrival permutation $\alpha$ returns allocation $M^{\prime\prime}$, which is not $IR$ since agent 5 receives item $e_{1}$.
In a first step, let me run Algorithm 7 on this instance in order to compute the departing agent excluded partition $\gamma(I)$. At $d_{1}$, $N_{<d_{1}}$ contains agents 1, 2, and 3. Therefore, $\{1\}$ and $\{2,3\}$ are added to the partition. At $d_{2}$ and $d_{3}$, nothing occurs since agents 2 and 3 already belong to the partition. At $d_{4}$, $N_{<d_{4}}$ contains the five agents, but agents 1, 2, and 3 already belong to the partition. Therefore, $\{4\}$ and $\{5\}$ are added to the partition, and Algorithm 7 halts, returning the departing agent excluded partition $\gamma(I)=\{\{1\},\{2,3\},\{4\},\{5\}\}$.
Let me now run Algorithm 6 with the departing agent excluded partition $\gamma(I)$ as input to compute the allocation for this instance. The TTC procedure is applied only once to a subset of more than one agent, which is $\{2,3\}$. As depicted in Figure 5, the TTC graph for the subset of agents $\{2,3\}$ contains a single cycle covering the two agents, and they swap their goods. The resulting allocation is $M$. Note that the same allocation is returned by Algorithm 3 using the ascending departure permutation $\alpha$, and therefore $M$ is $\mathcal{S}(I)$-$PO$.
@@
The idea under the departing agent excluded partition $\gamma$ is to punish agents, who force the assignment to take place because they leave earlier, by leaving them alone in the partition, with the consequence for Algorithm 6 that they keep their initial endowment without exchange. The following proposition shows more formally that no agent has an incentive to misreport her departure time when Algorithm 6 is used with the departing agent excluded partition $\gamma$.
@@
Note first that I know by Propositions 16 and 17 that Algorithm 6 using the departing agent excluded partition $\gamma$ is $WIC$. It remains to show that no agent has an incentive to misreport her departure time. Let $S$ denote the subset of agents for which the “if” clause of Algorithm 7 is true, i.e., agents that are not partitioned until their departure time.
@@
Consider first the case of an agent $i$ of $S$. Note first that agent $i$ belongs to a singleton in the departing agent excluded partition $\gamma$. Let me show that even if agent $i$ declares an earlier departure time $d^{\prime}_{i}$ such that $a_{i}<d^{\prime}_{i}<d_{i}$, then there is no other agent $j$ of $S$ such that $d^{\prime}_{i}<d_{j}<d_{i}$ holds. Indeed, since $d_{j}<d_{i}$, I know that agent $j$ is considered before agent $i$ during the “for” loop of Algorithm 6. In the last instruction of this “for” loop when agent $j$ is considered, the set of agents $N_{<d_{j}}$ is added to the set of partitioned agents $B$. Since agent $i$ is not part of $B$ at iteration $d_{i}$ of the “for” loop, it means that she does not belong to $N_{<d_{j}}$, or in other words $a_{i}>d_{j}$ holds. This implies, together with $a_{i}<d^{\prime}_{i}$, that $d_{j}<d^{\prime}_{i}$ holds.
@@
I now know that any agent of $S$ that leaves earlier than agent $i$ will still leave earlier than $i$ even after $i$ declares a new departure time $d^{\prime}_{i}$. This implies that agent $i$ will still pass the “if” test in the “for” loop of Algorithm 7 even if she changes her departure time to $d^{\prime}_{i}$ because no agent of $S$ leaving earlier will include agent $i$ in a partition. Therefore, agent $i$ will still be part of a singleton in the departing agent excluded partition $\gamma$ after misreporting her departure time and will keep her own item during Algorithm 6. Consequently, agent $i$ has no incentive to misreport her departure time.
@@
Consider now the case of an agent $i$ that is not part of $S$. Note first that the fact that $i$ is not part of $S$ means that there exists an agent $j$ of $S$ such that $d_{j}<d_{i}$ and $i\in N_{<d_{j}}$ hold. I assume that $j$ is the agent of $S$ with the lowest rank in the ascending departure permutation $\delta$ for which these inequalities are true. In other words, the subset of agents containing agent $i$ is selected by Algorithm 7 during the iteration of the “for” loop corresponding to $d_{j}$. By applying the same reasoning as above, it is easy to check that there is no agent $k$ of $S$ such that $d_{k}<d_{j}$ and $d^{\prime}_{i}<d_{k}$. Therefore, the only misreport of agent $i$ that could change the outcome of Algorithm 7 is a departure time such that $d_{j}>d^{\prime}_{i}$ holds. In that case, agent $i$ will belong to the set of agents for which the “if” condition of Algorithm 7 is true, and therefore agent $i$ will be part of a singleton in the departing agent excluded partition $\gamma$, and will receive her own item in Algorithm 6. Since Algorithm 6 is $IR$, agent $i$ has no incentive to misreport her departure time since she cannot receive a less preferred item than her own. ∎
@@ 7.2 Partition Functions Based on Schedules
Let me now examine if $a$-$IC$ can be achieved by Algorithm 6 using another partition function. To do so, I need to introduce an additional element $\xi=\{\xi_{0},\xi_{1},\ldots,\xi_{k}\}$, that I call scheduling, which is a subset of non-overlapping time intervals. More formally, I assume that for any $j\in\{0,1,\ldots,k\}$, $\xi_{j}\subseteq\mathcal{T}$ holds, and for any $j^{\prime}\neq j$, $\xi_{j}\cap\xi_{j^{\prime}}=\emptyset$ holds. I assume that each interval is contiguous, potentially with the exception of the first one $\xi_{0}$, which covers all time intervals not contained in the other intervals of the scheduling. This scheduling will be used to initialize the partition by grouping agents whose departure times belong to the same time interval of $\xi$.
The construction of the scheduled departure partition $\theta$, which is described in Algorithm 8, depends on both the arrival and departure times of the agents, and the input scheduling $\xi$. Intuitively, all agents whose departure times belong to the same time interval of $\xi$ and whose arrival times are no later than the earliest departure time of one of these agents are grouped together, whereas the other agents are left alone in the partition. The first interval of $\xi$, denoted $\xi_{0}$, plays a special role since all the agents whose departure times belong to this interval are left alone in the partition. Note that this first interval may be left empty. To simplify notation, I denote by $\theta$ the scheduled departure partition resulting from Algorithm 8, without specifying the scheduling used. This is not a big issue here since the properties proved later will not depend on the scheduling used as input. The following proposition shows that the partition function resulting from Algorithm 8 is static online compatible.
@@
Input: Instance $I$, and scheduling $\xi=\{\xi_{0},\xi_{1},\ldots\xi_{k}\}$.
1:  $A\leftarrow\{0\}$. {Set of times intervals of $\xi$ already considered.}
@@
7:        Let $S$ be the subset of agents of $I_{<d_{\delta(i)}}$ whose departure times belong to $\xi_{j}$.
@@
The scheduled departure partition $\theta$ is $SOC$.
@@
I need to show that for any instance $I$ and for any agent $i$, $\theta_{i}(I)=\theta_{i}(I_{<d_{i}})$ holds. Let $i^{\prime}$ denote the rank of agent $i$ in the ascending departure permutation $\delta$, i.e., such that $\delta(i^{\prime})=i$ holds. Note that for any position $j\leq i^{\prime}$, agents whose arrival times are no later than $d_{\delta(j)}$ are all contained in $I_{<d_{i}}$. Therefore, the first $i^{\prime}$ iterations of the “for” loop are the same in $I$ and $I_{<d_{i}}$. Thus, the subset of agents containing agent $i$ in the scheduled departure partition $\theta$ is the same for both instances $I$ and $I_{<d_{i}}$. ∎
@@
Consider once again instance $I$ described in the left part of Figure 4, and scheduling $\xi$ such that $[a_{3},d_{2}]\subseteq\xi_{1}$ and $[d_{3},d_{5}]\subseteq\xi_{2}$ hold. The timeline is reproduced in Figure 6 to provide a graphical representation of the scheduling $\xi$. Note that $\xi_{1}$ contains the departure time of agents 1 and 2, and $\xi_{2}$ contains the departure time of agents 3, 4, and 5. Let me run Algorithm 8 to compute the scheduled departure partition $\theta(I)$. At $d_{1}$, both agents 1 and 2 are in $N_{<d_{1}}$. Therefore, $\{1,2\}$ is added to the partition. At $d_{2}$, nothing happens since agent 2 already belongs to the partition. At $d_{3}$, both agents 3 and 4 are in $N_{<d_{3}}$. Therefore, $\{3,4\}$ is added to the partition. At $d_{4}$, nothing happens since agent 4 is already part of the partition. At $d_{5}$, the interval $\xi_{2}$ that contains $d_{5}$ has already been considered. Therefore, agent 5 remains alone, and $\{5\}$ is added to the partition. Finally, the algorithm halts and returns the scheduled departure partition $\theta(I)=\{\{1,2\},\{3,4\},\{5\}\}$.
@@
Let me now run Algorithm 6 using the scheduled departure partition $\theta(I)$ as input to compute the allocation. The TTC procedure is applied independently to $\{1,2\}$ and $\{3,4\}$. The TTC-graphs during the first iteration of the TTC procedure are represented in Figure 7. In the left part of the figure, we can see the TTC-graph for the subset of agents $\{1,2\}$, which contains a single cycle, the self-loop involving agent 2. Therefore, agent 2 keeps her good, and agent 1, who remains alone during the second iteration, also keeps her good. The TTC-graph shown in the right part of the figure corresponds to the subset of agents $\{3,4\}$. This TTC-graph contains a single cycle that includes both agents, who swap their goods. The resulting matching $M^{\prime}$ is depicted in the right part of Figure 4.
@@
Note first that I know by Propositions 16 and 20 that Algorithm 6 using the scheduled departure partition $\theta$ as input is $WIC$. It remains to show that no agent has an incentive to misreport her arrival time. Note first that since the departure time should be revealed truthfully by each agent, no agent can change the time interval of $\xi$ that is associated with her. Note also that if an agent’s departure time belongs to $\xi_{0}$, then whatever her reported arrival time is, she will be alone in the scheduled departure partition $\theta$ and keep her own item during Algorithm 6. Therefore, she has no incentive to misreport her true arrival time. I can now focus on agents whose arrival time does not belong to $\xi_{0}$.
@@
Consider the following instance $I$ with 2 agents such that $a_{1}<a_{2}<\xi_{1}^{-}<d_{2}<\xi_{1}^{+}<d_{1}$ hold, where $\xi_{1}^{-}$ and $\xi_{1}^{+}$ denote the starting and ending times, respectively, of $\xi_{1}$ [^8] Such starting and ending times should exist since I have assumed that time intervals are contiguous, with the exception of $\xi_{0}$.. Furthermore, the departure time $d_{1}$ belongs to interval $\xi_{0}$. It is easy to check that $\theta_{1}(I)=\{1\}$. Assume that agent 1 prefers $e_{2}$ to $e_{1}$, and agent 2 prefers $e_{1}$ to $e_{2}$. If agent 1 reveals her true departure time $d_{1}$, then Algorithm 6 using the scheduled departure partition $\theta$ will assign her item $e_{1}$. On the other hand, if she misreports her departure time and reports $d^{\prime}_{1}$ such that $\xi_{1}^{-}<d^{\prime}_{1}<\xi_{1}^{+}$, then assuming that $I^{\prime}$ denotes the copy of $I$ where agent 1 reveals $d^{\prime}_{1}$ instead of $d_{1}$, it is easy to check that $\theta_{1}(I^{\prime})=\{1,2\}$. Therefore, according to the preferences of agents 1 and 2, item $e_{2}$ is assigned by Algorithm 6 to agent 1. Since $e_{2}$ is the most favorite item of agent 1, she has an incentive to misreport her departure time. ∎
If the scheduling contains a single interval $\xi_{0}=T$, then the scheduled departure partition $\theta$ is the partition where each agent is alone, and Algorithm 6 will assign to each agent her own item. This algorithm is obviously $SIC$ since its outcome does not change with the preferences and the arrival and departure time of the agents. However, this algorithm is not really interesting since no exchanges among agents are performed, and the allocation is not improved. By combining the properties of Algorithms 7 and 8, I can create a less trivial partition function that renders Algorithm 6 $SIC$. To do so, I can use Algorithm 8 as a baseline, and modify line 8 by the assignment $\theta\leftarrow\theta\cup\{\{\delta(i)\}\}\cup\{S\setminus\{\delta(i)\}\}$, which is consistent with line 5 of Algorithm 7. Furthermore, I also restrict the input scheduling to the one containing two intervals such that $\xi_{0}=\emptyset$ and $\xi_{1}=T$. Let $\zeta$ denote the resulting partition function that I call earliest departure partition. In other words, the earliest departure partition $\zeta$ will group together all the agents of $I_{<\delta(1)}$, with the exception of agent 1, and all the other agents stay alone. Note that the earliest departure partition $\zeta$ can be easily shown to be $SOC$ since it results from a very similar algorithm to Algorithms 7 and 8.

### v2

@@ 7 Top Trading Cycle Based Algorithms
Let usfirst show that Algorithm 6, denoted $\mathcal{A}$ in this proof, is online when the partition function at use is $PP$. Consider an instance $I$ and an agent $i$. Let $I^{\prime}$ denote the restricted instance $I_{<d_{i}}$. We need to show that $\mathcal{A}_{i}(I)=\mathcal{A}_{i}(I^{\prime})$. Since $a_{i}<d_{j}$ holds for any agent $j$ of $I^{\prime}$, it follows that $I_{<d_{j}}=I^{\prime}_{<d_{j}}$. This implies that $\mathcal{P}_{j}(I_{<d_{j}})=\mathcal{P}_{j}(I^{\prime}_{<d_{j}})$ holds for any agent $j$ in $I^{\prime}$. Therefore, the partitions of agents in $I^{\prime}$ are the same in $\mathcal{P}(I)$ and $\mathcal{P}(I^{\prime})$, and Algorithm 6, applied to both $I$ and $I^{\prime}$, executes the same iterations up to the departure time $d_{i}$ of agent $i$. This implies that $\mathcal{A}_{i}(I)=\mathcal{A}_{i}(I^{\prime})$.
Let us now show that it is $WIC$. It is well known that Algorithm 5 is incentive compatible. During Algorithm 6, multiple instances of Algorithm 5 are applied, but each agent participates in exactly a single one of them. Furthermore, we assumed that the partition function used to select the agents participating in each instance of Algorithm 5 does not depend on the agents’ preferences. Therefore, no agent has an incentive to misreport her preferences, and Algorithm 6 is $WIC$. ∎
@@ 7.1 The Departing Agent Excluded Partition
We explore now if a stronger notion of incentive compatibility can be achieved with particular partition functions. Let us introduce the departing agent excluded partition $\gamma$, whose construction is described in Algorithm 7. Intuitively, every time that an unpartitioned agent leaves the market, two new subsets are added to the partition. The first is a singleton containing the agent leaving the market, and the second is the remaining unpartitioned agents that are present in the market. The following proposition shows that this partition function is progress-preserving.
@@
The departing agent excluded partition $\gamma$ is $PP$
@@
We need to show that for any instance $I$ and for any agent $i$, $\gamma_{i}(I)=\gamma_{i}(I_{<d_{i}})$ holds. Let $i^{\prime}$ denote the rank of agent $i$ in the ascending departure permutation $\delta$, i.e., such that $\delta(i^{\prime})=i$ holds. Note that for any position $j\leq i^{\prime}$, agents in $N_{<d_{\delta(j)}}$ are all contained in $I_{<d_{i}}$ since $d_{i}\geq d_{\delta(j)}$. Therefore, the first $i^{\prime}$ iterations of the “for” loop of Algorithm 7 are the same for $I$ and $I_{<d_{i}}$. Consequently, the subset of agents containing agent $i$ in the departing agent excluded partition $\gamma$ is the same for both instances $I$ and $I_{<d_{i}}$. ∎
@@
Consider the instance $I$ depicted in the left part of Figure 4, and the three allocations described in the right part of the same figure. It is easy to verify that both Algorithms 1 and 2 using either the ascending departure permutation $\delta$ or the ascending arrival permutation $\alpha$ return allocation $M^{\prime\prime}$, which is not $IR$ since agent 5 receives item $e_{1}$.
In a first step, let us run Algorithm 7 on this instance in order to compute the departing agent excluded partition $\gamma(I)$. At $d_{1}$, $N_{<d_{1}}$ contains agents 1, 2, and 3. Therefore, $\{1\}$ and $\{2,3\}$ are added to the partition. At $d_{2}$ and $d_{3}$, nothing occurs since agents 2 and 3 already belong to the partition. At $d_{4}$, $N_{<d_{4}}$ contains the five agents, but agents 1, 2, and 3 already belong to the partition. Therefore, $\{4\}$ and $\{5\}$ are added to the partition, and Algorithm 7 halts, returning the departing agent excluded partition $\gamma(I)=\{\{1\},\{2,3\},\{4\},\{5\}\}$.
Let us now run Algorithm 6 with the departing agent excluded partition $\gamma(I)$ as input to compute the allocation for this instance. The TTC procedure is applied only once to a subset of more than one agent, which is $\{2,3\}$. As depicted in Figure 5, the TTC graph for the subset of agents $\{2,3\}$ contains a single cycle covering the two agents, and they swap their goods. The resulting allocation, denoted by $M$, is illustrated in Figure 4. Note that the same allocation is returned by Algorithm 3 using the ascending departure permutation $\delta$, and therefore $M$ is $\mathcal{S}(I)$-$PO$.
@@
The idea under the departing agent excluded partition $\gamma$ is to “punish” agents, who force the assignment to take place because they leave earlier, by leaving them alone in the partition, with the consequence for Algorithm 6 that they keep their initial endowment without exchange. The following proposition shows more formally that no agent has an incentive to misreport her departure time when Algorithm 6 is used with the departing agent excluded partition $\gamma$.
@@
Note first that we know by Propositions 16 and 17 that Algorithm 6 using the departing agent excluded partition $\gamma$ is $WIC$. It remains to show that no agent has an incentive to misreport her departure time. Let $S$ denote the subset of agents for which the “if” clause of Algorithm 7 is true, i.e., agents that are not partitioned until their departure time.
@@
Consider first the case of an agent $i$ of $S$. Note first that agent $i$ belongs to a singleton in the departing agent excluded partition $\gamma$. Let us show that even if agent $i$ declares an earlier departure time $d^{\prime}_{i}$ such that $a_{i}<d^{\prime}_{i}<d_{i}$, then there is no other agent $j$ of $S$ such that $d^{\prime}_{i}<d_{j}<d_{i}$ holds. Indeed, since $d_{j}<d_{i}$, we know that agent $j$ is considered before agent $i$ during the “for” loop of Algorithm 6. In the last instruction of this “for” loop when agent $j$ is considered, the set of agents $N_{<d_{j}}$ is added to the set of partitioned agents $B$. Since agent $i$ is not part of $B$ at iteration $d_{i}$ of the “for” loop, it means that she does not belong to $N_{<d_{j}}$, or in other words $a_{i}>d_{j}$ holds. This implies, together with $a_{i}<d^{\prime}_{i}$, that $d_{j}<d^{\prime}_{i}$ holds.
@@
We now know that any agent of $S$ that leaves earlier than agent $i$ will still leave earlier than $i$ even after $i$ declares a new departure time $d^{\prime}_{i}$. This implies that agent $i$ will still pass the “if” test in the “for” loop of Algorithm 7 even if she changes her departure time to $d^{\prime}_{i}$ because no agent of $S$ leaving earlier will include agent $i$ in a partition. Therefore, agent $i$ will still be part of a singleton in the departing agent excluded partition $\gamma$ after misreporting her departure time and will keep her own item during Algorithm 6. Consequently, agent $i$ has no incentive to misreport her departure time.
@@
Consider now the case of an agent $i$ that is not part of $S$. Note first that the fact that $i$ is not part of $S$ means that there exists an agent $j$ of $S$ such that $d_{j}<d_{i}$ and $i\in N_{<d_{j}}$ hold. We assume that $j$ is the agent of $S$ with the lowest rank in the ascending departure permutation $\delta$ for which these inequalities are true. In other words, the subset of agents containing agent $i$ is selected by Algorithm 7 during the iteration of the “for” loop corresponding to $d_{j}$. By applying the same reasoning as above, it is easy to check that there is no agent $k$ of $S$ such that $d_{k}<d_{j}$ and $d^{\prime}_{i}<d_{k}$. Therefore, the only misreport of agent $i$ that could change the outcome of Algorithm 7 is a departure time such that $d_{j}>d^{\prime}_{i}$ holds. In that case, agent $i$ will belong to the set of agents for which the “if” condition of Algorithm 7 is true, and therefore agent $i$ will be part of a singleton in the departing agent excluded partition $\gamma$, and will receive her own item in Algorithm 6. Since Algorithm 6 is $IR$, agent $i$ has no incentive to misreport her departure time since she cannot receive a less preferred item than her own. ∎
@@ 7.2 Partition Functions Based on Schedules
Let us now examine if $a$-$IC$ can be achieved by Algorithm 6 using another partition function. To do so, we need to introduce an additional element $\xi=\{\xi_{0},\xi_{1},\ldots,\xi_{k}\}$, that we call scheduling, which is a subset of non-overlapping time intervals. More formally, we assume that for any $j\in\{0,1,\ldots,k\}$, $\xi_{j}\subseteq\mathcal{T}$ holds, and for any $j^{\prime}\neq j$, $\xi_{j}\cap\xi_{j^{\prime}}=\emptyset$ holds. We assume that each interval is contiguous, potentially with the exception of the first one $\xi_{0}$, which covers all time intervals not contained in the other intervals of the scheduling. This scheduling will be used to initialize the partition by grouping agents whose departure times belong to the same time interval of $\xi$.
The construction of the scheduled departure partition $\theta$, which is described in Algorithm 8, depends on both the arrival and departure times of the agents, and the input scheduling $\xi$. Intuitively, all agents whose departure times belong to the same time interval of $\xi$ and whose arrival times are no later than the earliest departure time of one of these agents are grouped together, whereas the other agents are left alone in the partition. The first interval of $\xi$, denoted $\xi_{0}$, plays a special role since all the agents whose departure times belong to this interval are left alone in the partition. Note that this first interval may be left empty. To simplify notation, we denote by $\theta$ the scheduled departure partition resulting from Algorithm 8, without specifying the scheduling used. This is not a big issue here since the properties proved later will not depend on the scheduling used as input. The following proposition shows that the partition function resulting from Algorithm 8 is progress-preserving.
@@
Input: Instance $I=\{(i,e_{i},a_{i},d_{i},\succ_{i})\}_{i\in N}$, and scheduling $\xi=\{\xi_{0},\xi_{1},\ldots\xi_{k}\}$.
1:  $A\leftarrow\{0\}$. {Set of time interval indices from $\xi$ that have already been considered.}
@@
7:        Let $S$ be the subset of agents of $N_{<d_{\delta(i)}}$ whose departure times belong to $\xi_{j}$.
@@
The scheduled departure partition $\theta$ is $PP$.
@@
We need to show that for any instance $I$ and for any agent $i$, $\theta_{i}(I)=\theta_{i}(I_{<d_{i}})$ holds. Let $i^{\prime}$ denote the rank of agent $i$ in the ascending departure permutation $\delta$, i.e., such that $\delta(i^{\prime})=i$ holds. Note that for any position $j\leq i^{\prime}$, agents whose arrival times are no later than $d_{\delta(j)}$ are all contained in $I_{<d_{i}}$. Therefore, the first $i^{\prime}$ iterations of the “for” loop are the same in $I$ and $I_{<d_{i}}$. Thus, the subset of agents containing agent $i$ in the scheduled departure partition $\theta$ is the same for both instances $I$ and $I_{<d_{i}}$. ∎
@@
Consider once again instance $I$ described in the left part of Figure 4, and scheduling $\xi$ such that $[a_{3},d_{2}]\subseteq\xi_{1}$ and $[d_{3},d_{5}]\subseteq\xi_{2}$ hold. The timeline is reproduced in Figure 6 to provide a graphical representation of the scheduling $\xi$. Note that $\xi_{1}$ contains the departure time of agents 1 and 2, and $\xi_{2}$ contains the departure time of agents 3, 4, and 5. Let us run Algorithm 8 to compute the scheduled departure partition $\theta(I)$. At $d_{1}$, both agents 1 and 2 are in $N_{<d_{1}}$. Therefore, $\{1,2\}$ is added to the partition. At $d_{2}$, nothing happens since agent 2 already belongs to the partition. At $d_{3}$, both agents 3 and 4 are in $N_{<d_{3}}$. Therefore, $\{3,4\}$ is added to the partition. At $d_{4}$, nothing happens since agent 4 is already part of the partition. At $d_{5}$, the interval $\xi_{2}$ that contains $d_{5}$ has already been considered. Therefore, agent 5 remains alone, and $\{5\}$ is added to the partition. Finally, the algorithm halts and returns the scheduled departure partition $\theta(I)=\{\{1,2\},\{3,4\},\{5\}\}$.
@@
Let us now run Algorithm 6 using the scheduled departure partition $\theta(I)$ as input to compute the allocation. The TTC procedure is applied independently to $\{1,2\}$ and $\{3,4\}$. The TTC-graphs during the first iteration of the TTC procedure are represented in Figure 7. In the left part of the figure, we can see the TTC-graph for the subset of agents $\{1,2\}$, which contains a single cycle, the self-loop involving agent 2. Therefore, agent 2 keeps her good, and agent 1, who remains alone during the second iteration, also keeps her good. The TTC-graph shown in the right part of the figure corresponds to the subset of agents $\{3,4\}$. This TTC-graph contains a single cycle that includes both agents, who swap their goods. The resulting matching $M^{\prime}$ is depicted in the right part of Figure 4.
@@
Note first that we know by Propositions 16 and 20 that Algorithm 6 using the scheduled departure partition $\theta$ as input is $WIC$. It remains to show that no agent has an incentive to misreport her arrival time. Note first that since the departure time should be revealed truthfully by each agent, no agent can change the time interval of $\xi$ that is associated with her. Note also that if an agent’s departure time belongs to $\xi_{0}$, then whatever her reported arrival time is, she will be alone in the scheduled departure partition $\theta$ and keep her own item during Algorithm 6. Therefore, she has no incentive to misreport her true arrival time. We can now focus on agents whose arrival time does not belong to $\xi_{0}$.
@@
Consider the following instance $I$ with 2 agents such that $a_{1}<a_{2}<\xi_{1}^{-}<d_{2}<\xi_{1}^{+}<d_{1}$ hold, where $\xi_{1}^{-}$ and $\xi_{1}^{+}$ denote the starting and ending times, respectively, of $\xi_{1}$ [^13] Such starting and ending times should exist since we have assumed that time intervals are contiguous, with the exception of $\xi_{0}$.. Furthermore, the departure time $d_{1}$ belongs to interval $\xi_{0}$. It is easy to check that $\theta_{1}(I)=\{1\}$. Assume that agent 1 prefers $e_{2}$ to $e_{1}$, and agent 2 prefers $e_{1}$ to $e_{2}$. If agent 1 reveals her true departure time $d_{1}$, then Algorithm 6 using the scheduled departure partition $\theta$ will assign her item $e_{1}$. On the other hand, if she misreports her departure time and reports $d^{\prime}_{1}$ such that $\xi_{1}^{-}<d^{\prime}_{1}<\xi_{1}^{+}$, then assuming that $I^{\prime}$ denotes the copy of $I$ where agent 1 reveals $d^{\prime}_{1}$ instead of $d_{1}$, it is easy to check that $\theta_{1}(I^{\prime})=\{1,2\}$. Therefore, according to the preferences of agents 1 and 2, item $e_{2}$ is assigned by Algorithm 6 to agent 1. Since $e_{2}$ is the most favorite item of agent 1, she has an incentive to misreport her departure time. ∎
7.3 The Threshold-Triggered Partition
We conclude this section by introducing a partition function that incentivizes agents to truthfully reveal both their arrival and departure times. First, note that if the scheduling consists of a single interval $\xi_{0}=T$, then the scheduled departure partition $\theta$ corresponds to the partition in which each agent forms a singleton set. In this case, Algorithm 6 simply assigns each agent her own item. This algorithm is trivially $SIC$ since its outcome does not depend on the agents’ preferences or their arrival and departure times. However, it is not particularly interesting, as no exchanges are performed and the resulting allocation is not improved.
By combining the properties of Algorithms 7 and 8, we construct a more meaningful partition function that still ensures Algorithm 6 is $SIC$. This function, denoted $\zeta$ and referred to as the threshold-triggered partition, is defined by Algorithm 9. The algorithm is equivalent to Algorithm 8, except that line 8 is replaced by the assignment $\theta\leftarrow\theta\cup\{\{\delta(i)\}\}\cup\{S\setminus\{\delta(i)\}\}$ (mirroring line 6 of Algorithm 7), and the input is restricted to a scheduling consisting of exactly two intervals, namely $\xi_{0}=[t^{-},\tau]$ and $\xi_{1}=[\tau,t^{+}]$ for a given threshold $\tau\in\mathcal{T}$ [^14] These two intervals are unambiguously determined by the threshold $\tau$, which therefore replaces in Algorithm 9 the scheduling as the input. Let $j$ denote the first agent whose departure time exceeds the threshold $\tau$. Then, the threshold-triggered partition $\zeta$ groups together all agents in $N_{<d_{j}}$, excluding agent $j$ and any agents who departed before $\tau$. All other agents are placed in singleton sets. Finally, note that $\zeta$ is clearly $PP$, as it is derived from a variant of the algorithms defining the departing agent excluded partition $\theta$ and the scheduled departure partition $\gamma$.
Algorithm 9 Construction of the threshold-triggered partition $\zeta$.
Input: Instance $I=\{(i,e_{i},a_{i},d_{i},\succ_{i})\}_{i\in N}$, and a threshold $t\in\mathcal{T}$.
1:  $thresholdPassed\leftarrow\mathtt{false}$. {Boolean that becomes true when $d_{\delta(i)}$ exceeds $t$.}
3:  for $i=1$ to $n$ do
4:     {Iteration occurring at $d_{\delta(i)}$.}
5:     if $\neg thresholdPassed$ and $d_{\delta(i)}\geq t$ then
6:        $thresholdPassed\leftarrow\mathtt{true}$.{Threshold exceeded for the first time.}
7:        $\theta\leftarrow\theta\cup\{\{\delta(i)\}\}\cup\{N_{<d_{\delta(i)}}\setminus(B\cup\{\delta(i)\})\}$.
8:        $B\leftarrow N_{<d_{\delta(i)}}$.
9:     else
10:        $\theta\leftarrow\theta\cup\{\{\delta(i)\}\}$.
11:        $B\leftarrow B\cup\{\delta(i)\}$.
12:     end if
13:  end for